OpenStudy (anonymous):

Please help me to prove Sin^2 -Cos^2 =1 .....

OpenStudy (anonymous):

Sin^2(x)-Cos^2(x) does not equal one.

OpenStudy (anonymous):

\[ \sin ^{2} - \cos ^{2} =1 \] Please Prove this....

OpenStudy (anonymous):

\[\sin ^2(x)-\cos^2(x)\neq1\]

OpenStudy (anonymous):

But I ve got this from a web site.... Here's the address.... Please go there and see whats wrong with it... .... http://www.cramster.com/practice-problems-p25-0-0-0-0-0-39167-0-cpi0-0.aspx

OpenStudy (anonymous):

Which number?

OpenStudy (anonymous):

In the Practice-3 Question set.... number 4

OpenStudy (anonymous):

\[\sin ^2(x)+\cos^2(x)=1\]\[\sin ^2(x)-\cos^2(x)=-\cos(2x)\]

OpenStudy (anonymous):

Oh.... Friutless can you please explain this farther......

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

Your reply for the Question I 've asked.....

OpenStudy (anonymous):

ok well this can be proved for sin(x)^2 + cos(x)^2 = 1 because this identity is true. you want to prove this? THE IDENTITY sin(x)^2 - cos(x)^2 = 1 IS NOT TRUE. DON'T THINK ABOUT WHAT THE WEBSITE SAYS. IT IS NOT TRUE. PERIOD. what would you like help proving?

OpenStudy (anonymous):

It is asking which one from the choices is correct. It is obviously not D.

OpenStudy (anonymous):

(i) is the correct answer.

OpenStudy (anonymous):

Ok prove this one then... \[\sin ^{8}A-\cos ^{8}A = (\sin ^{2}A-\cos ^{2}A) (1- 2 \sin ^{2}A \cos ^{2}A)\]

OpenStudy (anonymous):

apply the FOIL method to the right side (expand it) \[-\cos^2(A)+\sin^2(A)+2 \cos^4(A) \sin^2(A)-2 Cos^2(A) Sin^4(A)\]

OpenStudy (anonymous):

Replace the first two functions with an identity, and factor the rest.

OpenStudy (anonymous):

Hmmmm.... friutless in which grade you're studying... ?

OpenStudy (anonymous):

high school

OpenStudy (anonymous):

In which country...?

OpenStudy (anonymous):

usa

OpenStudy (anonymous):

Really.... then I think you have done something similar to the syllbus of London A/L isn't you?

OpenStudy (anonymous):

I am asking this from you because I am a student who has just started doing A/L s in Sri Lanka (somewhat similar to London A/L ). so I don't know the FOIL method..... So I can't understand you're reply.....

OpenStudy (anonymous):

Given (x-a)(y-b), applying FOIL yeilds xy-xb-ay+ab

OpenStudy (anonymous):

Do the same on the trigonometric expressions

OpenStudy (anonymous):

wow... I've got the right teacher here... thank you....

OpenStudy (anonymous):

did you get the left side from this method?

OpenStudy (anonymous):

\[Sin ^{8}A-\cos ^{8}A=( \sin ^{2}A-\cos ^{2}A)(\sin ^{6}A+\cos ^{6}A)+\sin ^{2}Acos ^{6}A-\cos ^{2}Asin ^{6}A\] Here's what I've done....... but I've stucked in it please try this also....

OpenStudy (anonymous):

Please help me......